# A variational approach to dimension-free self-normalized concentration

### v1

@@ 5 Bernstein and Bennett Inequalities
which, as long as the log-MGF term $\log\mathbb{E}_{k-1}\exp(\lambda\langle\theta,X\rangle)$ is finite, is easily seen to satisfy $\mathbb{E}_{t-1}M_{t}(\theta)=M_{t-1}(\theta)$ and is hence a nonnegative martingale. Depending on the assumptions we make on $(X_{t})$, we can bound the log-MFG by some function of $\mathbb{E}_{k-1}X_{k}X_{k}^{\intercal}$, which will give rise to a sub-$\psi$ process. The following lemma showcases this when the random vectors are bounded in each direction. The resulting process will provide our self-normalized Bennett’s inequality.
@@
Let $(X_{t})$ be a stream of random vectors in $\mathbb{R}^{d}$ with conditional mean 0 and $\sup_{\theta\in\mathbb{S}^{d-1}}\langle\theta,X_{t}\rangle\leq b$ almost surely for some constant $b>0$. Then $S_{t}=\sum_{j\leq t}X_{j}$ and $V_{t}=U_{0}+\sum_{j\leq t}\mathbb{E}_{j-1}X_{j}X_{j}^{\intercal}$ constitute a sub-$\psi_{P,b}$ process for any PSD matrix $U_{0}$.
@@
Using these two results in conjunction with Theorem 4.1 we obtain the following line crossing inequalities. Figure 4 plots $\psi^{-1}_{P,1}$ versus $\psi^{-1}_{G,1}$ to give a sense of the constraints on $\lambda$ in the following result. It also plots $\psi_{G,1}$ versus $\psi_{P,1}$ to give some sense of how (5.3) scales compared to (5.4).
@@
Let $(X_{t})$ be a sequence of random vectors with conditional mean 0. Suppose either that (i) $\sup_{\theta\in\mathbb{S}^{d-1}}\langle\theta,X_{t}\rangle\leq b$ for all $t\geq 1$ and some $b>0$ or (ii) that (5.2) holds. Let $S_{t}=\sum_{k\leq t}X_{k}$ and $V_{t}=U_{0}+\sum_{k\leq t}\mathbb{E}_{k-1}[X_{k}X_{k}^{\intercal}]$ for some positive definite matrix $U_{0}$. Set $\rho=\|U_{0}\|$. Fix any $\delta\in(0,1)$. If (i) holds, then for any $\lambda\in[\psi_{P,b}^{-1}(1/\gamma_{\min}(U_{0})),\infty)$, with probability $1-\delta$, for any stopping time $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{\sqrt{\rho}}{\lambda g_{\psi_{P,b},\tau}(\lambda)}\left(\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det U_{0}}\right)+1+\log(1/\delta)\right)+\frac{g_{\psi_{P,b},\tau}(\lambda)\psi_{P,b}(\lambda)}{\lambda\sqrt{\rho}},$ (5.3)
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{\sqrt{\rho}}{\lambda g_{\psi_{G,c},\tau}(\lambda)}\left(\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det U_{0}}\right)+1+\log(1/\delta)\right)+\frac{g_{\psi_{G,c},\tau}(\lambda)\psi_{G,c}(\lambda)}{\lambda\sqrt{\rho}}.$ (5.4)
Since the process is sub-gamma when $(X_{t})$ satisfy the Bernstein condition in (5.2), we can in that case apply Theorem 4.3 to obtain a stitched boundary. We omit the statement here for brevity, but we will provide a similar result in the next section when studying an empirical Bernstein bound.
@@ 6 An Empirical Bernstein Inequality
In this section we provide a self-normalized “empirical Bernstein” result. Empirical Bernstein bounds are useful because they do not require a priori knowledge of the variance in order to be instantiated. Instead, they adapt to the unknown variance. Such bounds have been studied in scalar random variables [45, 58, 34], vector-valued random variables [37, 42, 43, 14], u-statistics [50], and matrices [57]. As far as we are aware, Whitehouse et al. [60] gave the first empirical Bernstein result for self-normalized, vector-valued processes. Ours is thus the first dimension-free, self-normalized, empirical Bernstein inequality.
Let $(X_{t})$ be a stream of random vectors with conditional mean 0. Suppose that $\|X_{t}\|\leq 1/2$ (one can replace 1/2 with any constant and rescale the bound accordingly). Recall that $\psi_{E,1}(\lambda)=-\log(1-\lambda)-\lambda$. Denoting $\widehat{\mu}_{t}=\frac{1}{t}\sum_{j\leq t}X_{j}$, the process defined by
$N_{t}(\theta)=\prod_{k\leq t}\exp\left\{\lambda\langle\theta,X_{k}\rangle-\psi_{E,1}(\lambda)\langle\theta,X_{k}-\widehat{\mu}_{k-1}\rangle^{2}\right\},\quad N_{0}(\theta)=1$ (6.1)
is a nonnegative supermartingale (see Chugg et al. [14, Lemma E.1] for an explicit proof). It is a multivariate extension of a supermartingale used by Howard et al. [34], which they use to prove an empirical Bernstein bound for scalar random variables. This supermartingale is, in turn, an extension of an inequality of Fan et al. [21], where the mean is replaced by the key term $\widehat{\mu}_{k-1}$ in the empirical variance. This reduces the variance when the mean of $X_{k}$ is nonzero and delivers a certain asymptotic sharpness property of the resulting confidence set widths in scalar, vector, and matrix settings [58, 42, 57].
@@
$S_{t}=\sum_{k\leq t}X_{k},\quad V_{t}=\rho I_{d}+\sum_{k\leq t}(X_{k}-\widehat{\mu}_{k-1})(X_{k}-\widehat{\mu}_{k-1})^{\intercal},$ (6.2)
for any fixed $\rho\geq 0$. (The term $\rho I_{d}$ is not in (6.1), but adding to $V_{t}$ only lowers the value of the process.) Since $\psi_{E,1}$ is super-Gaussian, we may apply Theorem 4.1 to obtain the following result.
@@
Let $(X_{t})$ be a stream of random vectors such that $\|X_{t}\|_{2}\leq 1/2$ for all $t\geq 1$. Let $S_{t}=\sum_{k=1}^{t}X_{k}$ and $V_{t}=\rho I_{d}+\sum_{k=1}^{t}(X_{k}-\widehat{\mu}_{k-1})(X_{k}-\widehat{\mu}_{k-1})^{\intercal}$ for any $\rho>1$. Let $g_{t}(z)=g_{\psi_{E,1},t}(z)$ for $g_{\psi,t}$ as in (4.3). Then, for any $\delta\in(0,1)$ and any $\lambda\in[\psi_{E,1}^{-1}(1/\rho),1)$, with probability $1-\delta$, for any stopping time $\tau$,
@@
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{\sqrt{\rho}}{\lambda g_{\tau}(\lambda)}\left(\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det\rho I_{d}}\right)+1+\log(1/\delta)\right)+\frac{g_{\tau}(\lambda)\psi_{E,1}(\lambda)}{\lambda\sqrt{\rho}}.$ (6.3)
In Corollary 6.1, we need to constrain $\rho$ to be larger than 1 to ensure that a a valid domain exists for $\lambda$. To elaborate, in Theorem 4.1 we require that $\lambda\in[\psi^{-1}(1/\gamma_{\min}(U_{0}),\lambda_{\max})$. Here, $\psi^{-1}(1/\gamma_{\min(U_{0})})=\psi^{-1}_{E,1}(1/\rho)$ and $\lambda_{\max}=1$. To have $\psi^{-1}_{E,1}(1/\rho)<1$ requires that $\rho>1$.
A skeptical reader might eye the term $\det(\rho I_{d})$ in (6.3) with suspicion. This term depends on $d$, after all! However, $\rho I_{d}$ can be replaced with any $U_{0}$ such that $\gamma_{\min}(U_{0})>1$, which would make the bound truly dimension-free. We set $U_{0}=\rho I_{d}$ only for easier comparison to previous work.
@@
In order to remove the dependence on $\lambda$ in Corollary 6.1, we again turn to stitching. Since stitching with $\psi=\psi_{E,1}$ is difficult to do directly, we instead upper bound $\psi_{E,c}$ with $\psi_{G,c}$ using the following lemma, and then apply Theorem 4.3.
@@
As illustrated by Figure 4(a), for small $\lambda$, $\psi_{E,1}$ and $\psi_{G,1}$ are nearly identical. They separate more as $\lambda$ grows (note that $\lambda_{\max}=1$). Using Lemma 6.2 and applying Theorem 4.3 with $(S_{t},V_{t})$ as in (6.2) gives the following result. As in Corollary 6.1, we require that $\rho>1$ in order to ensure that $\psi_{G,1}^{-1}(1/\rho),1)$ is non-empty.
@@
Suppose $\|X_{t}\|_{2}\leq 1/2$. Let $V_{t}=\rho I_{d}+\sum_{k=1}^{t}(X_{k}-\widehat{\mu}_{k-1})(X_{k}-\widehat{\mu}_{k-1})^{\intercal}$ for any $\rho>1$. Fix $\delta\in(0,1/\sqrt{2}]$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{V_{\tau}^{-1}}\leq\frac{D_{\tau}+1.60\sqrt{D_{\tau}}+\max\left\{\frac{1+\sqrt{1+2\rho}}{2\rho},\sqrt{\frac{D_{\tau}}{2}}\right\}}{H_{\tau}},$ (6.4)
@@
$D_{\tau}=\frac{1}{2}\log\left(\frac{\det V_{\tau}}{\det(\rho I_{d})}\right)+1.5+2\log(\log_{2}(\det V_{\tau})+1)+\log(1/\delta),$ (6.5)
@@
$H_{\tau}=0\vee\left(\sqrt{1-\frac{\rho}{\gamma_{\min}(V_{t})}}-\sqrt{\frac{2\rho}{\gamma_{\min}(V_{t})(1+\sqrt{3})}}\right).$ (6.6)
7 Summary

### v2

@@ 5 Self-normalized Bernstein and Bennett Inequalities
which, as long as the log-MGF term $\log\mathbb{E}_{k-1}\exp(\lambda\langle\theta,X\rangle)$ is finite, is easily seen to satisfy $\mathbb{E}_{t-1}M_{t}(\theta)=M_{t-1}(\theta)$ and is hence a nonnegative martingale. (We use the superscript $B$ to signal Bennett/Bernstein). Depending on the assumptions we make on $(X_{t})$, we can bound the log-MFG by some function of $\mathbb{E}_{k-1}X_{k}X_{k}^{\intercal}$, which will give rise to a sub-$\psi$ process. The following lemma showcases this when the random vectors are bounded in each direction. The resulting process will provide our self-normalized Bennett’s inequality.
@@
Let $(X_{t})$ be a stream of random vectors in $\mathbb{R}^{d}$ with conditional mean 0 obeying $\sup_{\theta\in\mathbb{S}^{d-1}}\langle\theta,X_{t}\rangle\leq b$ almost surely for some constant $b>0$. Then $S_{t}=\sum_{j\leq t}X_{j}$ and $V_{t}=U_{0}+\sum_{j\leq t}\mathbb{E}_{j-1}X_{j}X_{j}^{\intercal}$ constitute a sub-$\psi_{P,b}$ process for any PSD matrix $U_{0}$.
@@
Using these two results in conjunction with Theorem 4.1 we obtain the following line crossing inequalities. Figure 3 plots $\psi^{-1}_{P,1}$ versus $\psi^{-1}_{G,1}$ to give a sense of the constraints on $\lambda$ in the following result. It also plots $\psi_{G,1}$ versus $\psi_{P,1}$ to indicate how (5.4) scales compared to (5.5).
@@
Let $(X_{t})$ be a sequence of random vectors with conditional mean 0. Let $U_{0}$ be positive definite. Suppose either that (i) $\sup_{\theta\in\mathbb{S}^{d-1}}\langle\theta,X_{t}\rangle\leq b$ for all $t\geq 1$ and some $b>0$ or (ii) that (5.2) holds. Let $S_{t}=\sum_{k\leq t}X_{k}$ and $V_{t}=\sum_{k\leq t}\mathbb{E}_{k-1}[X_{k}X_{k}^{\intercal}]$. Fix any $\delta\in(0,1)$ and let
$D_{t}=\frac{1}{2}\log\left(\frac{\det(V_{\tau}+U_{0})}{\det U_{0}}\right)+1+\log(1/\delta).$ (5.3)
If (i) holds, then for any $\lambda\in[\psi_{P,b}^{-1}(1/\gamma_{\min}(U_{0})),\infty)$, with probability $1-\delta$, for any stopping time $\tau$,
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{D_{t}}{\lambda g_{\psi_{P,b},\tau}(\lambda)}+\frac{g_{\psi_{P,b},\tau}(\lambda)\psi_{P,b}(\lambda)}{\lambda},$ (5.4)
@@
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{D_{t}}{\lambda g_{\psi_{G,c},\tau}(\lambda)}+\frac{g_{\psi_{G,c},\tau}(\lambda)\psi_{G,c}(\lambda)}{\lambda}.$ (5.5)
Since the process is sub-gamma when $(X_{t})$ satisfies the Bernstein condition in (5.2), we can in that case apply Theorem 4.4 to obtain the stitched boundary in (4.5). In particular, by carefully applying the line-crossing inequality above with different values of $\lambda$ at different times, we are able to apply (5.5) with $\lambda$ as roughly $\sqrt{D_{t}}/(1+c\sqrt{D_{t}})$, up to constant multipliers on $D_{t}$. This gives $\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}$ scaling like $cD_{\tau}+\sqrt{D_{t}}$ with high probability. This combination of linear and sublinear terms is what we expect of a Bernstein bound.
@@ 6 An Empirical Bernstein Inequality
In this section we provide a self-normalized “empirical Bernstein” result. Empirical Bernstein bounds are useful because they do not require a priori knowledge of the variance in order to be instantiated. Instead, they adapt to the unknown variance. Such bounds have been studied in scalar random variables [47, 60, 35, 14], vector-valued random variables [38, 45, 44, 16], u-statistics [52], and matrices [59, 14]. As far as we are aware, Whitehouse et al. [63] gave the first empirical Bernstein result for self-normalized, vector-valued processes. Ours is thus the first dimension-free, self-normalized, empirical Bernstein inequality.
Let $(X_{t})$ be a stream of random vectors with conditional mean $\mu$. Suppose that $\|X_{t}\|\leq 1/2$ (one can replace 1/2 with any constant and rescale the bound accordingly). Recall that $\psi_{E,1}(\lambda)=-\log(1-\lambda)-\lambda$. For any adapted sequence $(\widehat{\mu}_{t})_{t\geq 0}$, [^6] That is, $\widehat{\mu}_{t}$ is $\mathcal{F}_{t}$-measurable. the process defined by
$N_{t}(\theta)=\prod_{k\leq t}\exp\left\{\lambda\langle\theta,X_{k}-\mu\rangle-\psi_{E,1}(\lambda)\langle\theta,X_{k}-\widehat{\mu}_{k-1}\rangle^{2}\right\},\quad N_{0}(\theta)=1.$ (6.1)
is a nonnegative supermartingale (see Chugg et al. [16, Lemma E.1] for an explicit proof). It is a multivariate extension of a supermartingale used by Howard et al. [35], which they use to prove an empirical Bernstein bound for scalar random variables. This supermartingale is, in turn, an extension of an inequality of Fan et al. [23], where the mean is replaced by the key term $\widehat{\mu}_{k-1}$ in the empirical variance. This reduces the variance when the mean of $X_{k}$ is nonzero and delivers a certain asymptotic sharpness property of the resulting confidence set widths in scalar, vector, and matrix settings [60, 45, 59].
@@
$S_{t}=\sum_{k\leq t}(X_{k}-\mu),\quad V_{t}=\Gamma+\sum_{k\leq t}(X_{k}-\widehat{\mu}_{k-1})(X_{k}-\widehat{\mu}_{k-1})^{\intercal},$ (6.2)
and $\Gamma$ is any fixed PSD matrix (could be the zero matrix). Since $\psi_{E,1}$ is super-Gaussian, we may apply Theorem 4.1 to obtain the following result.
@@
Let $(X_{t})$ be a stream of random vectors such that $\|X_{t}\|_{2}\leq 1/2$ for all $t\geq 1$. Let $U_{0}$ be positive definite and $V_{t}$ be as above. Let $g_{t}(z)=g_{\psi_{E,1},t}(z)$ for $g_{\psi,t}$ as in (4.3). Then, for any $\delta\in(0,1)$ and any $\lambda\in[\psi_{E,1}^{-1}(1/\gamma_{\min}(U_{0})),1)$, with probability $1-\delta$, for any stopping time $\tau$,
$\bigg\|\sum_{k\leq\tau}(X_{k}-\mu)\bigg\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{1}{\lambda g_{\tau}(\lambda)}\left(\frac{1}{2}\log\left(\frac{\det(V_{\tau}+U_{0})}{\det U_{0}}\right)+1+\log(1/\delta)\right)+\frac{g_{\tau}(\lambda)\psi_{E,1}(\lambda)}{\lambda}.$
The upper bound of 1 on $\lambda$ comes from the fact that $\lambda_{\max}=1$ for $\psi_{E,1}$. Note that $\psi^{-1}(1/\kappa)<1$ for all $\kappa>0$ so there is no constraint on the minimum eigenvalue of $U_{0}$ other than positivity.
We added the PSD matrix $\Gamma$ to the process $(V_{t})$ in Corollary 6.1 because we can often obtain better bounds in practice by splitting $U_{0}=\Gamma+U_{0}^{\prime}$. To elaborate, suppose we want to bound $\|S_{t}\|_{(V_{t}+U_{0})^{-1}}$ for some fixed $U_{0}$ using Corollary 6.1. If the $X_{k}$ are mostly pointing in one direction, then $V_{t}$ will be low-rank, in which case $\alpha_{t}=\alpha(U_{0},U_{0}+V_{t})\approx 1$ and $g_{t}(\lambda)\approx 0$, blowing up the bound. Thus, applying the bound as stated can lead to vacuous results, at least until $\gamma_{\min}(V_{t})$ grows sufficiently large. However, we can also apply the bound with $U_{0}^{\prime}=(1-\epsilon)U_{0}$ and $\Gamma=\epsilon U_{0}$ for any $0<\epsilon<1$, which ensures that $\alpha_{t}$ is bounded away from 1. Of course, this is bounding the same process since $V_{t}+\Gamma+U_{0}^{\prime}=V_{t}+U_{0}$ by construction. This leads to much better results in practice.
Figure 4 employs this stabilization trick and compares Corollary 6.1 to the self-normalized empirical Bernstein bound of Whitehouse et al. [63, Theorem 4.1]. Translated to our setting, this latter bounds reads: For all $\delta\in(0,1)$, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{t}\|_{(V_{\tau}+U_{0})^{-1}}\leq 2\sqrt{\gamma_{\min}(V_{\tau}+U_{0})}\cdot(\psi_{E,1}^{*})^{-1}\left(\frac{2L(V_{\tau})}{\gamma_{\min}(V_{\tau}+U_{0})}\right),$ (6.3)
where
@@
$L(V_{t})=\log\left(\frac{2}{\delta}\right)+\log\left(h\left(\log_{2}\left(\frac{\gamma_{\max}(V_{t}+U_{0})}{\gamma_{\min}(U_{0})}\right)\right)\right)+\log\left(2\sqrt{\kappa_{t}}\cdot N_{d-1}\left(\frac{1/4}{\sqrt{\kappa_{t}}}\right)\right),$
and $\kappa_{t}=\kappa(V_{t}+U_{0})$ is the condition number of $V_{t}+U_{0}$. We have simplified the bound somewhat by fixing certain parameters that are left as variables in the original. Here $N_{d-1}(\epsilon)$ is the $\epsilon$-covering number of the unit sphere $\mathbb{S}^{d-1}$, i.e., the minimum size of an $\epsilon$-cover of $\mathbb{S}^{d-1}$. This is where the dimension-dependence comes from, as $N_{d-1}(\epsilon)=\Omega(\epsilon^{-(d-1)})$.
When comparing our bound with (6.3) we take $(\widehat{\mu}_{t})$ to be the sequence of sample means, i.e., $\widehat{\mu}_{t}=t^{-1}\sum_{k\leq t}X_{k}$. The results echo those in Section 4: our bound beats the condition number-based bound of Whitehouse et al. insofar as $\log\det(V_{t})\lesssim d\log(\kappa(V_{t}))$, where $\kappa(V_{t})$ is the condition number of $V_{t}$. As the determinant grows relative to the condition number, our bound becomes looser than (6.3). In practice, if we do not know how the determinant will grow, we might use a union bound to take advantage of both bounds.
(a) k=2,d=20k=2,\;d=20
(b) k=5,d=20k=5,\;d=20
(c) k=10,d=20k=10,\;d=20
Figure 4: Comparison of our empirical Bernstein bound, Corollary 6.1 (blue line) , with the empirical Bernstein bound of Whitehouse et al. [63] (dotted line). Here each $X_{t}$ is generated to have random uniform noise in $k$ directions, so $V_{t+1}=V_{t}+(X_{t+1}-\widehat{\mu}_{t})(X_{t+1}-\widehat{\mu}_{t})^{\top}$ acts roughly as a rank-$k$ update. As $k$ grows, the determinant grows relative to the condition number, thus our determinant-based bound deteriorates with respect to the condition number bound. Simulation details can be found in Appendix C.
@@
In order to remove the dependence on $\lambda$ in Corollary 6.1, we may again turn to stitching. Since stitching with $\psi=\psi_{E,1}$ is difficult to do directly, we instead upper bound $\psi_{E,c}$ with $\psi_{G,c}$ using the following lemma, and then apply Theorem 4.4.
Here we provide the details on the stitched empirical Bernstein bound.
@@
As illustrated by Figure 3(a), for small $\lambda$, $\psi_{E,1}$ and $\psi_{G,1}$ are nearly identical. They separate more as $\lambda$ grows (note that $\lambda_{\max}=1$). Using Lemma 6.2 and applying Theorem 4.4 with $(S_{t},V_{t})$ as in (6.2) gives the following result.
@@
Suppose $\|X_{t}\|_{2}\leq 1/2$. Let $V_{t}=\sum_{k=1}^{t}(X_{k}-\widehat{\mu}_{k-1})(X_{k}-\widehat{\mu}_{k-1})^{\intercal}$ and let $U_{0}$ be positive definite with $\upsilon=\gamma_{\min}(U_{0})$. Fix $\delta\in(0,1/e]$. Then, with probability $1-\delta$, for all stopping times $\tau$,
$\|S_{\tau}\|_{(V_{\tau}+U_{0})^{-1}}\leq\frac{D_{\tau}+1.51\sqrt{D_{\tau}}+0.74\max\left\{\frac{1+\sqrt{1+2\upsilon}}{2\upsilon},\sqrt{\frac{D_{\tau}}{2}}\right\}}{H_{\tau}},$ (6.4)
@@
$D_{\tau}=\frac{1}{2}\log\left(\frac{\det(V_{\tau}+U_{0})}{\det U_{0}}\right)+1.5+2\log(\log_{2}(\det(V_{\tau}+U_{0}))+1)+\log(1/\delta),$ (6.5)
@@
$H_{\tau}=0\vee\left(0.42-0.86\sqrt{\frac{\gamma_{\max}(U_{0})}{\gamma_{\min}(V_{\tau}+U_{0})}}\right).$ (6.6)
The precise numerical values in the bounds come from instantiating the various constants in Theorem 4.4 with $c=1$ and simplifying. As with the stitching results in Section 4.2, Corollary 6.3 is mostly of theoretical interest to understand the rate of our bounds. In practice we suggest using Corollary 6.1.
7 Conclusion